OpenStudy (alyssajobug):

Solve the system of linear equations and check any solution algebraically. x + y + z = 2 -x + 3y + 2z = 8 4x + y = 4

OpenStudy (anonymous):

do you know how to use the elimination method? probably, what we can do is to eliminate z variable first using the 1st and 2nd equation then use the resulting equation apply another elimination method to solve for the variable x or y ^_^ so first, what can you do to the 1st and 2nd equation to eliminate z?

OpenStudy (alyssajobug):

So could we multiply the 1st equation by 2 and then subtract it from the 2nd equasion?

OpenStudy (anonymous):

yes exactly :)

OpenStudy (anonymous):

i'm trying to solve it... just tell me what equation you'll get after you applied what you said and let's compare it to mine, sounds good?

OpenStudy (alyssajobug):

ok, so after multiplying and subtracting you would get x + y + z = 2 3x - y = 8 4x + y = 4 right?

OpenStudy (anonymous):

wait, you should do what you said: \(\color{blue}{\text{Originally Posted by}}\) @alyssajobug So could we multiply the 1st equation by 2 and then subtract it from the 2nd equasion? \(\color{blue}{\text{End of Quote}}\) multiply 1st equation by 2 : so it will be 2x+2y+2z=2 then subtract this by 2nd equation

OpenStudy (anonymous):

you should end up with one equation (leave equation 3 for now)

OpenStudy (alyssajobug):

So x + y + z = 2 3x - y =4 4x + y = 4 ?

OpenStudy (anonymous):